OpenStudy (anonymous):

first correct answer will get a medal!! what is the slope of the line parallel to the line whose equation is 5x+3y=15

OpenStudy (anonymous):

do you know what the slope of 5x + 3y = 15 is ?

OpenStudy (anonymous):

put the equation in y = mx + b form, and the number in the m position is the slope. A parallel line will have the same slope

OpenStudy (anonymous):

i forget to put this up a.)-5/3 b.)-3/5 c.)5/3 d.)3/5

OpenStudy (anonymous):

can you put this equation into y = mx + b form ?

OpenStudy (anonymous):

ok....I am in a generous mood 5x + 3y = 15 ---- subtract 5x from both sides 3y = -5x + 15 --- divide by 3 y = -5/3x + 15/3 y = -5/3x + 5 the number in the m position (the slope) is -5/3 A parallel line will have the same slope.

OpenStudy (nirmalnema):

slopes of two parallel lines are equal.. so you can find a slope of a given line by using the formula.. -b/a a=coeff... of x and b = coef. of y so its slope is - 5/3

OpenStudy (anonymous):

wow...thats easier nirmalnema.

OpenStudy (nirmalnema):

yup.. :)

OpenStudy (anonymous):

smarty pants....lol

OpenStudy (anonymous):

yeah I was setting it up but im kinda really new to thing and thank you for your help

OpenStudy (anonymous):

anytime :)

OpenStudy (nirmalnema):

lol....!!